OpenStudy (anonymous):

A rectangle has perimeter 84cm and length 35cm. What is it's width?

OpenStudy (anonymous):

91?

OpenStudy (anonymous):

Perimeter of a rectangle = length + length + width + width So, plugging in the numbers we're given, 84 = 35 + 35 + w + w 84 = 70 + 2w 14 = 2w 7 = w(idth)!

OpenStudy (anonymous):

It's not adding

OpenStudy (anonymous):



OpenStudy (anonymous):

@onemillion6 do not give full answers